what constant force p is required to bring a 50kg body which starts from rest to a speed of 10m/s in 7m neglect friction

Grade:12th pass

2 Answers

Ramreddy IIT Patna
9 years ago
use v2 -u2 =2*a*s given u=0 and s= 7 m and v=10 m/s
100= 2*a*7
a= 100/14
Force required(p)= ma = 50*100/14 = 357 N
….….….….….….….…..............................................
Aziz Alam IIT Roorkee
9 years ago
a=v2/2s
=>a=100/14=7.14 m/s2

F=ma=50*7.14=357.14 N


------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------